# Covariant dynamics from static spherically symmetric geometries

### v2

@@ Acknowledgements.
Appendix A Solving the dynamics
Under the PG gauge, the evolution equation $\dot{E}^{I}=0$ leads to
$0=$ $-\frac{2N}{x}\frac{\partial{H_{\rm eff}}}{\partial s_{3}}+2xN^{x},$ (A1)
$0=$ $-N\frac{\partial{H_{\rm eff}}}{\partial s_{2}}+\partial_{x}(xN^{x}).$
Due to Eq. (1), we have
$\frac{\partial{H_{\rm eff}}}{\partial s_{2}}=$ $-\partial_{x}(\partial_{s_{2}}M_{\rm eff})-2x\partial_{M_{\rm eff}}\mathcal{R}\partial_{s_{2}}M_{\rm eff},$
$\frac{\partial{H_{\rm eff}}}{\partial s_{3}}=$ $-x\partial_{s_{2}}M_{\rm eff},$
where the diffeomorphism constraint $H_{x}=0$ is applied for the first equation. Thus, Eq. (A1) can be simplified to be
$0=$ $N\partial_{s_{2}}M_{\rm eff}+xN^{x},$
$0=$ $N\partial_{x}(\partial_{s_{2}}M_{\rm eff})+2Nx\partial_{M_{\rm eff}}\mathcal{R}\partial_{s_{2}}M_{\rm eff}+\partial_{x}(xN^{x}),$
which finally leads to Eq. (5). For $\dot{M}_{\rm eff}$, a straightforward calculation shows
$\dot{M}_{\rm eff}=-\frac{N\partial_{s_{2}}M_{\rm eff}}{E^{2}}\left(2E^{2}\mathcal{R}+\partial_{x}M_{\rm eff}\right)=0,$
where $E^{2}=\partial_{x}E^{1}/2$ as a result of the PG gauge and the first equation in Eq. (A1) are applied for the first equality, and the PG gauge and Eq. (4) are applied to get the last equality.
Appendix B Solving the covariance equation
Let us start from Eq. (6). By defining $Q=\partial_{s_{2}}M_{\rm eff}$, we can rewrite it as

### v3

@@ Acknowledgements.
Appendix —Equation (4) can be verified easily using the identity Zhang et al. (2024)
${H_{\rm eff}}-2\frac{\partial_{s_{2}}M_{\rm eff}}{\partial_{x}E^{1}}H_{x}=-\frac{2E^{2}}{\partial_{x}E^{1}}\left(\mathcal{R}\partial_{x}E^{1}+\partial_{x}M_{\rm eff}\right).$
Moreover, substituting this identity into the evolution equation $\partial_{t}M_{\rm eff}=\{M_{\rm eff},{H_{\rm eff}}[N]+H_{x}[N^{x}]\}$ and imposing $H_{x}=0$, we obtain
$\partial_{t}M_{\rm eff}+2N\mathcal{R}\partial_{s_{2}}M_{\rm eff}=N^{x}\partial_{x}M_{\rm eff}.$ (A1)
Additionally, with $H_{x}=0$ applied, the evolution equations for $E^{I}$ read
$\dot{E}^{1}$ $=2N\partial_{s_{2}}M_{\rm eff}+N^{x}\partial_{x}E^{1},$ (A2)
$\dot{E}^{2}$ $=\partial_{x}(N^{x}E^{2})+N\frac{2E^{2}}{\partial_{x}E^{1}}\partial_{x}(\partial_{s_{2}}M_{\rm eff})$
$+2NE^{2}(\partial_{M_{\rm eff}}\mathcal{R})\partial_{s_{2}}M_{\rm eff}.$
The first line gives a relation between $N$ and $N^{x}$. Inserting this relation into Eq. (A1) and using Eq. (4), we find $\partial_{t}M_{\rm eff}+\mathcal{R}\partial_{t}E^{1}=0$.
Let us impose a gauge fixing that specifies $E^{I}(t,x)$ ($I=1,2$). On any coordinate patch where $\partial_{s_{2}}M_{\rm eff}\neq 0$, the evolution equation (A2) determines $N$ and $N^{x}$, yielding Eqs. (5) and (6). On patches where $\partial_{s_{2}}M_{\rm eff}=0$, Eq. (A2) no longer fixes $N$. In this case, the evolution equation of $\partial_{s_{2}}M_{\rm eff}$ is needed. A direct calculation shows
$\partial_{x}N-N\left((\partial_{x}E^{1})\partial_{M_{\rm eff}}\mathcal{R}+\partial_{x}\ln(\frac{\partial_{x}E^{1}}{E^{2}})\right)$
$=$ $\frac{2(E^{2})^{2}}{\mu E^{1}\partial_{x}E^{1}}\left(\partial_{t}\left(\partial_{s_{2}}M_{\rm eff}\right)-\frac{N\partial_{x}\left[\left(\partial_{s_{2}}M_{\rm eff}\right)^{2}\right]}{\partial_{x}E^{1}}\right)$
$=$ $0,$
where the first equality is obtained by substituting Eqs. (3a) and (4) and by the vanishing of $\{\partial_{s_{2}}M_{\rm eff},H_{x}[N^{x}]\}=N^{x}\partial_{x}(\partial_{s_{2}}M_{\rm eff})$ on patches with $\partial_{s_{2}}M_{\rm eff}=0$. This result leads to Eq. (5) with $\partial_{x}C=0$.
Appendix —Let us start from Eq. (7). By defining $Q=\partial_{s_{2}}M_{\rm eff}$, we can rewrite it as